Reports 1-1 of 1 Clear search Modify search
MIF (General)
takafumi.ushiba - 22:46 Thursday 10 April 2025 (33347) Print this report
TF from PRCL displacement to DARM dispacement

As suggested from Komori-kun, I measured the TF from PRCL displacement to DARM displacement to evaluate the noise coupling from PRCL displacement to DARM displacement (fig1).
Unforetunately, IFO was down during the measurement but TF from 20Hz to 300Hz was measured.

Images attached to this report
Comments to this report:
takaaki.yokozawa - 8:18 Friday 11 April 2025 (33350) Print this report
I performed the noise projection based on the TF measured by this log
Images attached to this comment
takaaki.yokozawa - 8:41 Friday 11 April 2025 (33351) Print this report
I also performed the measurement for MICH.
But we need some more investigation for the noise projection
Images attached to this comment
takaaki.yokozawa - 9:52 Friday 11 April 2025 (33357) Print this report
As pointed out by Ushiba-san, the MICH displacement above 20 Hz would be sensing noise, so please ignore my projection of MICH.

As pointed out by Komori-san, we need tto multiply the effect of the OLG 1/(1+G) to the projection of PRCL, so I performed it.
Images attached to this comment
Search Help
×

Warning

×